# Enabling decision support over confidential data

### v1

@@ VIII Related Work
In recent years, numerous approaches have been proposed in the field of decision-making support systems. Haarmann et al. [14] propose executing DMN decision models on the Ethereum blockchain. However, blockchain-based approaches inherently rely on public ledgers, making them unsuitable for confidential data applications. In contrast, our TEE-based solution enables automated decision-making support securely and privately while maintaining the trust, transparency, and audibility required in multi-party processes. Moreover, in our solution, functions’ access policies ensure that only authorized participants can interact with the system, thus receiving decision-making support or confidential data. Zhang et al. [58] introduce a privacy-preserving decision-making system over blockchain for secure voting. In this paper, the authors proposed a low complexity Distributed Batch Key Generation and then designed new Two-Stage Voting scheme to reduce voting effort. This solution solves the privacy issues that arise when using blockchain technology, but it is tailored for decision-making voting only. On the contrary, SPARTA allows for user-definable decision logic. Other solutions, such as [13] and [12], allow for privacy-preserving clinical decisions for automated classification. These solutions resort to homomorphic encryption to maintain data confidentiality, resulting in significant performance overhead due to encryption. They resort to a Naïve Bayes classifier and a Gaussian kernel-based SVM as the decision support system. In contrast, our solution uses AES-GCM encryption, introducing low overhead to the overall performance and allowing for specifying the decision logic for the decision-making support. Liu et al. [59] present HPCS, a Hybrid Privacy-Preserving Clinical Decision Support System used for health status monitoring and disease prediction. Here, a server securely monitors patients’ conditions in real-time. The new symptoms can be sent to the cloud server running a neural network for prediction in a privacy-preserving way. Different from SPARTA, the authors use the Paillier Cryptosystem with Threshold Decryption to guarantee data privacy, adding high overhead on the overall procedure. Moreover, HPCS limits the decision logic to the sole predictions, while we allow for user-definable logic. Xie et al. [60] present TEBDS, a data-sharing system combining TEEs with blockchain. Their design leverages a TEE for off-chain storage and a consortium blockchain for access control in IoT contexts, with a centralized key management center (KMC). In contrast, our approach relies on a public blockchain only for storing IPNS names pointing to IPFS content-addressed data while delegating all operations to the TEE. This design eliminates the need for a centralized KMC and enables confidential decision-making over encrypted data.
IX Conclusion

### v2

@@ 8 Related Work
In recent years, numerous approaches have been proposed in the field of decision-making support systems. Haarmann et al. [haarmann2018dmn] propose executing DMN decision models on the Ethereum blockchain. However, blockchain-based approaches inherently rely on public ledgers, making them unsuitable for confidential data applications. In contrast, our TEE-based solution enables automated decision-making support securely and privately while maintaining the trust, transparency, and audibility required in multi-party processes. Moreover, in our solution, functions’ access policies ensure that only authorized participants can interact with the system, thus receiving decision-making support or confidential data. Zhang et al. [Zhang-privacydecisions] introduce a privacy-preserving decision-making system over blockchain for secure voting. In this paper, the authors proposed a low complexity Distributed Batch Key Generation and then designed new Two-Stage Voting scheme to reduce voting effort. This solution solves the privacy issues that arise when using blockchain technology, but it is tailored for decision-making voting only. On the contrary, SPARTA allows for user-definable decision logic. Other solutions, such as [Liu] and [Rahulamathavan], allow for privacy-preserving clinical decisions for automated classification. These solutions resort to homomorphic encryption to maintain data confidentiality, resulting in significant performance overhead due to encryption. They resort to a Naïve Bayes classifier and a Gaussian kernel-based SVM as the decision support system. In contrast, our solution uses AES-GCM encryption, introducing low overhead to the overall performance and allowing for specifying the decision logic for the decision-making support. Liu et al. [LIU2018825] present HPCS, a Hybrid Privacy-Preserving Clinical Decision Support System used for health status monitoring and disease prediction. Here, a server securely monitors patients’ conditions in real-time. The new symptoms can be sent to the cloud server running a neural network for prediction in a privacy-preserving way. Different from SPARTA, the authors use the Paillier Cryptosystem with Threshold Decryption to guarantee data privacy, adding high overhead on the overall procedure. Moreover, HPCS limits the decision logic to the sole predictions, while we allow for user-definable logic. Xie et al. [TEBDS] present TEBDS, a data-sharing system combining TEEs with blockchain. Their design leverages a TEE for off-chain storage and a consortium blockchain for access control in IoT contexts, with a centralized key management center (KMC). In contrast, our approach relies on a public blockchain only for storing IPNS names pointing to IPFS content-addressed data while delegating all operations to the TEE. This design eliminates the need for a centralized KMC and enables confidential decision-making over encrypted data.
9 Conclusion
@@
Appendix A Running example decisions
Table VII: Decision model for the vaccination centers’ restock of Sect. 3
When And And And Then
CurrentVaccineStockLevel MaxStorageCapacity PopulationServed VaccinationProgress Restock
number number number number string
$\leq$ MaxStorageCapacity $\cdot$ 0.1 $\geq$1000 - - “Immediate”
$\leq$ MaxStorageCapacity $\cdot$ 0.25 $\geq$1000 >50000 <10000 “Needed soon”
$\leq$ MaxStorageCapacity $\cdot$ 0.25 [500..1000[ >20000 <5000 “Needed”
$\leq$ MaxStorageCapacity $\cdot$ 0.5 $\geq$1000 >1000 >10000 “Medium priority”
$\leq$ MaxStorageCapacity $\cdot$ 0.5 [500..1000[ [5000..30000] >20000 “Lower priority”
$\leq$ MaxStorageCapacity $\cdot$ 0.25 $\geq$1000 >50000 >45000 “No need”
Table VIII: Decision model for the carriers selection of Sect. 3
When And And And And And And And Then
CoverageArea NumberOfTrucks NumberOfPlanes OperationalCost PerKm RefrigerationCapacity Temperature MonitoringSystems AnnualDeliveries ISOCompliant Carrier
number number number number boolean boolean number boolean string
>2000 >10 $\geq$1 <5 true true >45000 true “Rapid Logistics Inc.”
$\geq$2000 $\geq$10 >1 [2..3] true false $\geq$15000 true “PrimeWay Haulage”
$\geq$2000 <10 >1 [1..2[ true false $\geq$20000 true “Atlas Cargo Express”
[1000..2000[ $\geq$5 >1 $\leq$5 true false $\geq$15000 true “SkyBridge Freight”
[500..1000[ $\geq$3 >1 $\leq$3.5 true false $\geq$25000 true “Global Freight Solutions”
[500..1000[ $\geq$3 >1 >3.5 true false $\geq$15000 true “Evergreen Transport Co.”
<500 $\geq$2 0 <2 false false $\geq$30000 true “Horizon Cargo Services”
<500 1 0 <2 true false $\geq$10000 true “Velocity Transport Ltd.”
<500 $\geq$5 1 <3 false true $\geq$30000 true “BlueLine Logistics”
- - - - - - - false “Not Eligible”
Listing 5: Decision function object derived from Tab. VII
1func Restock(inputs) {
2 if inputs["CurrentVaccineStockLevel"] <= inputs["MaxStorageCapacity"] * 0.1 and inputs["MaxStorageCapacity"] >= 1000 { return "Immediate" }
3 if inputs["CurrentVaccineStockLevel"] <= inputs["MaxStorageCapacity"] * 0.25 and inputs["MaxStorageCapacity"] >= 1000 and inputs["PopulationServed"] > 50000 and inputs["VaccinationProgress"] < 10000 { return "Needed soon" }
4 if inputs["CurrentVaccineStockLevel"] <= inputs["MaxStorageCapacity"] * 0.25 and inputs["MaxStorageCapacity"] >= 500 and inputs["MaxStorageCapacity"] < 1000 and inputs["PopulationServed"] > 20000 and inputs["VaccinationProgress"] < 5000 { return "Needed" }
5 if inputs["CurrentVaccineStockLevel"] <= inputs["MaxStorageCapacity"] * 0.5 and inputs["MaxStorageCapacity"] >= 1000 and inputs["PopulationServed"] > 1000 and inputs["VaccinationProgress"] > 10000 { return "Medium priority" }
6 if inputs["CurrentVaccineStockLevel"] <= inputs["MaxStorageCapacity"] * 0.5 and inputs["MaxStorageCapacity"] >= 500 and inputs["MaxStorageCapacity"] < 1000 and inputs["PopulationServed"] >= 5000 and inputs["PopulationServed"] <= 30000 and inputs["VaccinationProgress"] > 20000 { return "Lower priority" }
7 if inputs["CurrentVaccineStockLevel"] <= inputs["MaxStorageCapacity"] * 0.25 and inputs["MaxStorageCapacity"] >= 1000 and inputs["PopulationServed"] > 50000 and inputs["VaccinationProgress"] > 45000 { return "No need" }
8 return "No matching decision found" }
Listing 6: Decision function object derived from Tab. VIII
1func Carrier(inputs) {
2 if inputs["CoverageArea"] > 2000 and inputs["NumberOfTrucks"] > 10 and inputs["NumberOfPlanes"] >= 1 and inputs["OperationalCostPerKm"] < 5 and inputs["RefrigerationCapacity"] == true and inputs["TemperatureMonitoringSystems"] == true and inputs["AnnualDeliveries"] > 45000 and inputs["ISOCompliant"] == true { return "Rapid Logistics Inc." }
3 if inputs["CoverageArea"] >= 2000 and inputs["NumberOfTrucks"] >= 10 and inputs["NumberOfPlanes"] > 1 and inputs["OperationalCostPerKm"] >= 2 and inputs["OperationalCostPerKm"] <= 3 and inputs["RefrigerationCapacity"] == true and inputs["TemperatureMonitoringSystems"] == false and inputs["AnnualDeliveries"] >= 15000 and inputs["ISOCompliant"] == true { return "PrimeWay Haulage" }
4 if inputs["CoverageArea"] >= 2000 and inputs["NumberOfTrucks"] < 10 and inputs["NumberOfPlanes"] > 1 and inputs["OperationalCostPerKm"] >= 1 and inputs["OperationalCostPerKm"] < 2 and inputs["RefrigerationCapacity"] == true and inputs["TemperatureMonitoringSystems"] == false and inputs["AnnualDeliveries"] >= 20000 and inputs["ISOCompliant"] == true { return "Atlas Cargo Express" }
5 if inputs["CoverageArea"] >= 1000 and inputs["CoverageArea"] < 2000 and inputs["NumberOfTrucks"] >= 5 and inputs["NumberOfPlanes"] > 1 and inputs["OperationalCostPerKm"] <= 5 and inputs["RefrigerationCapacity"] == true and inputs["TemperatureMonitoringSystems"] == false and inputs["AnnualDeliveries"] >= 15000 and inputs["ISOCompliant"] == true { return "SkyBridge Freight" }
6 if inputs["CoverageArea"] >= 500 and inputs["CoverageArea"] < 1000 and inputs["NumberOfTrucks"] >= 3 and inputs["NumberOfPlanes"] > 1 and inputs["OperationalCostPerKm"] <= 3.5 and inputs["RefrigerationCapacity"] == true and inputs["TemperatureMonitoringSystems"] == false and inputs["AnnualDeliveries"] >= 25000 and inputs["ISOCompliant"] == true { return "Global Freight Solutions" }
7 if inputs["CoverageArea"] >= 500 and inputs["CoverageArea"] < 1000 and inputs["NumberOfTrucks"] >= 3 and inputs["NumberOfPlanes"] > 1 and inputs["OperationalCostPerKm"] > 3.5 and inputs["RefrigerationCapacity"] == true and inputs["TemperatureMonitoringSystems"] == false and inputs["AnnualDeliveries"] >= 15000 and inputs["ISOCompliant"] == true { return "Evergreen Transport Co." }
8 if inputs["CoverageArea"] < 500 and inputs["NumberOfTrucks"] >= 2 and inputs["NumberOfPlanes"] == 0 and inputs["OperationalCostPerKm"] < 2 and inputs["RefrigerationCapacity"] == false and inputs["TemperatureMonitoringSystems"] == false and inputs["AnnualDeliveries"] >= 30000 and inputs["ISOCompliant"] == true { return "Horizon Cargo Services" }
9 if inputs["CoverageArea"] < 500 and inputs["NumberOfTrucks"] == 1 and inputs["NumberOfPlanes"] == 0 and inputs["OperationalCostPerKm"] < 2 and inputs["RefrigerationCapacity"] == true and inputs["TemperatureMonitoringSystems"] == false and inputs["AnnualDeliveries"] >= 10000 and inputs["ISOCompliant"] == true { return "Velocity Transport Ltd." }
10 if inputs["CoverageArea"] < 500 and inputs["NumberOfTrucks"] >= 5 and inputs["NumberOfPlanes"] == 1 and inputs["OperationalCostPerKm"] < 3 and inputs["RefrigerationCapacity"] == false and inputs["TemperatureMonitoringSystems"] == true and inputs["AnnualDeliveries"] >= 30000 and inputs["ISOCompliant"] == true { return "BlueLine Logistics" }
11 if inputs["ISOCompliant"] == false { return "Not Eligible" }
12 return "No matching decision found" }
In the following, Tab. VII and Tab. VIII present the DMN tables containing the decision logic for the two remaining decisions of Sect. 3. The medical hub can request decision support for the restocking of each vaccination center and for selecting the carrier responsible for each delivery. Listing 5 and List. 6 show the two resulting DeFObjs corresponding to Tab. VII and Tab. VIII, respectively.